OpenStudy (anonymous):

Write three fractions that are equivalent to fifteen-twentieths. Show your wo

OpenStudy (anonymous):

\[\frac{ 15 }{ 20}\times2=\frac{ 30 }{ 40 } \div5=\frac{ 6 }{ 8 }\div2=\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

There you go!!!!